# Spike No More: Stabilizing the Pre-training of Large Language Models

### v2

@@ Appendix A Hyper-Parameters
Table 3 shows that hyper-parameters used in our experiments on LLMs. In addition to experiments described in Section 4, this table also indicates the hyper-parameters of the model with 13B parameters that we evaluated in Appendix E.
Appendix B Related Work
Stability
To stabilize trainings of Transformer-based neural language models, there have been various discussions on the architecture (Xiong et al., 2020; Liu et al., 2020; Takase et al., 2023; Zeng et al., 2023; Zhai et al., 2023), initialization method (Nguyen & Salazar, 2019; Zhang et al., 2019b; Huang et al., 2020; Wang et al., 2022), training strategy (Zhang et al., 2022; Li et al., 2022), and loss function (Chowdhery et al., 2022; Wortsman et al., 2023).
Xiong et al. (2020) theoretically analyzed gradient scales of each part in Transformers, and indicated that the Pre-LN Transformer is more stable than the Post-LN Transformer, that is the original Transformer architecture (Vaswani et al., 2017). Since the Pre-LN Transformer is more stable than the Post-LN Transformer theoretically and empirically, recent studies mainly have used the Pre-LN Transformer to construct an LLM. We also assume using the Pre-LN Transformer in the analysis on the training dynamics in this paper.
To stabilize the LLM pre-training, Le Scao et al. (2022) applied the layer normalization to the embedding layer. Zeng et al. (2023) used shrink embedding gradient technique (Ding et al., 2021). In this study, we proved that the layer normalization to the embedding layer prevents the exploding gradients around layer normalizations in internal layers when we use the scaled initialization, which is a widely used initialization method for LLMs (Nguyen & Salazar, 2019; Shoeybi et al., 2020), and thus, it stabilizes the pre-training.
For the initialization methods, Nguyen & Salazar (2019) proposed a strategy to initialize parameters of Transformers with small values to stabilize their training. Zhang et al. (2019b) and Huang et al. (2020) indicated that we can remove layer normalizations in Transformers if we use their proposed initialization methods. Wang et al. (2022) adjusted initial parameter scales based on the number of layers to stabilize the Post-LN Transformer. In this study, we indicated that an initialization method affects the LLM pre-training dynamics. In particular, we indicated that the widely used initialization method (Shoeybi et al., 2020) prevents the explosion due to the shortcut but causes the explosion due to layer normalizations. Moreover, we proved that we can prevent the explosion due to layer normalizations by making the variance of embeddings close to $1$.
Efficiency
As shown in Table 2, our modification enables the pre-training with a relatively larger learning rate, and can achieve better performance. Thus, this study can be regarded as on the efficiency of LLM pre-training because our modification can construct a better LLM with a given budget. Strubell et al. (2019) and Schwartz et al. (2019) reported that recent neural methods require substantial computational costs, and thus, they argued that we have to explore a cost-efficient approach.
Rajbhandari et al. (2020) proposed ZeRO that reduces memory redundancies during the multi GPU training without increasing communication volume. Dao et al. (2022) focused on GPU memory reads/writes, and proposed FlashAttention that accelerates the speed of attention mechanisms in Transformers. To reduce the number of computations in the attention mechanism, Shazeer (2019) proposed the multi-query attention that shares one key and value across all of the attention heads in each layer. Takase & Kiyono (2023) explored several parameter sharing strategies, and indicated that parameter sharing across some layers can achieve comparable performance to the vanilla model with a small number of parameters. Moreover, several studies have explored a better construction way with a limited budget (Izsak et al., 2021; Takase & Kiyono, 2021). We believe that we can take advantage of their findings to make our LLMs more efficient.
Appendix C Wang and Komatsuzaki Initialization
Black et al. (2022) and Biderman et al. (2023) initialized $W_{2}$ and $W_{O}$ with the normal distribution $\mathcal{N}(0,2/(N\sqrt{d}))$ instead of using $\mathcal{N}(0,\sigma/\sqrt{2N})$. We compare these two standard deviations: $2/(N\sqrt{d})$ and $(\sigma/\sqrt{2N})$ in this section. As described in Section 2.2, we use $\sigma=\sqrt{2/5d}$, and thus, $\sigma/\sqrt{2N}=1/\sqrt{5Nd}$.
Let $\upsilon$ be $1/\sqrt{5Nd}$, then $\upsilon\sqrt{20/N}=2/(N\sqrt{d})$. Therefore, if $N>20$, $\upsilon>2/(N\sqrt{d})$. In addition, the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ are much smaller than ones described in Section 3.2 in the case where we use the initialization method by Wang & Komatsuzaki (2021) and $\upsilon>2/(N\sqrt{d})$. As described in Appendix A, $N$ is larger than $20$ in this paper. Thus, we also face the exploding gradients by LNs if we use the initialization method by Wang & Komatsuzaki (2021).
Appendix D Comparison with Post-LN Transformer
@@
Appendix E Pre-training of the Model with 13B Parameters
(a) $\mathrm{lr}=3.0\times 10^{-4}$
(b) $\mathrm{lr}=1.0\times 10^{-4}$.
WikiText $\downarrow$ LAMBADA $\downarrow$
Vanilla N/A 15.12 N/A 6.50
Scaled Embed 14.47 15.25 5.97 6.53
Table 6 shows the perplexities of each configuration in evaluation data. This table indicates that we can achieve better performance when we use a larger learning rate in the same as in Section 5.1. In addition, the perplexities of Scaled Embed were comparable to ones of Vanilla when we used the small learning rate: $\mathrm{lr}=1.0\times 10^{-4}$. These results imply that our modification has no considerable risk in pre-training. Thus, we recommend making the standard deviation of each embedding close to $1$ to stabilize the pre-trainings of LLMs.

### v3

@@ Appendix A Hyper-Parameters
Appendix B Methods without Small Sub-Layers
Since we applied the widely used initialization method for LLMs in experiments in Section 5, all methods satisfy the condition on the small sub-layers. In this section, we empirically investigate the property of the method that violates the condition. We compare the Transformer initialized by the Xavier initialization (Glorot & Bengio, 2010) (Xavier Init), and the combination of Xavier Init and Scaled Embed (Xavier Init + Scaled Embed) with Scaled Embed. As described in Table 1, Xavier Init violates both conditions, and Xavier Init + Scaled Embed satisfies the only large shortcut condition. In the same manner as in Section 5, we trained models of 350M and 1.7B parameters with $\mathrm{lr}=5.0\times 10^{-4}$. We also used the hyper-parameters described in Table 4.
Figure 7 shows loss curves in validation data for 350M and 1.7B parameters in each method, and Figure 8 shows their gradient norms. These figures show that Xavier Init and Xavier Init + Scaled Embed faced loss and gradient spikes. In particular, the spikes appeared more frequently in Xavier Init, which violates both conditions, in comparison with Xavier Init + Scaled Embed. In contrast, Scaled Embed, which satisfies both conditions, avoided the gradient spike and prevented the loss spike problem. These results indicate that we have to satisfy both conditions: small sub-layers and large shortcut to prevent the loss spike problem. Moreover, Table 5 shows perplexities of each configuration in evaluation data. This table indicates that Scaled Embed achieved better performance than Xavier Init and Xavier Init + Scaled Embed that faced some spikes.
(a) 350M parameters.
(b) 1.7B parameters.
Figure 7: Loss curves of each method in validation data for the comparison to methods without small sub-layers.
(a) 350M parameters.
(b) 1.7B parameters.
Figure 8: Gradient norms of each method during the training for the comparison to methods without small sub-layers.
Table 5: Perplexities of each method.
Model WikiText $\downarrow$ LAMBADA $\downarrow$
350M parameters
Xavier Init 33.92 34.72
Xavier Init + Scaled Embed 30.50 26.55
Scaled Embed 29.86 24.37
1.7B parameters
Xavier Init 30.10 29.29
Xavier Init + Scaled Embed 23.16 15.49
Scaled Embed 21.29 12.53
Appendix C Pre-training of the Model with 13B Parameters
(a) $\mathrm{lr}=3.0\times 10^{-4}$
(b) $\mathrm{lr}=1.0\times 10^{-4}$.
WikiText $\downarrow$ LAMBADA $\downarrow$
Vanilla N/A 15.12 N/A 6.50
Scaled Embed 14.47 15.25 5.97 6.53
Table 6 shows the perplexities of each configuration in evaluation data. This table indicates that we can achieve better performance when we use a larger learning rate in the same as in Section 6.1. In addition, the perplexities of Scaled Embed were comparable to ones of Vanilla when we used the small learning rate: $\mathrm{lr}=1.0\times 10^{-4}$. These results imply that our modification has no considerable risk in pre-training. Thus, we have to satisfy large shortcut in addition to small sub-layers to stabilize the pre-trainings of LLMs.
Appendix D RMSNorm
Some recent LLMs use the RMSNorm (Zhang & Sennrich, 2019) instead of the LN in their Transformers (Touvron et al., 2023). We discuss such an architecture in this section. In the same as LN discussed in Section 3.1, we can obtain the Jacobian matrix of the RMSNorm with the following equation:
$\bigg{\|}\frac{\partial\mathrm{RMSNorm}(x)}{\partial x}\bigg{\|}_{2}=\frac{1}{\sigma_{x}}I$ (22)
Thus, the upper bound of the gradient norm is the same in LN if we use RMSNorm.
Figure 10 shows the loss values and gradient norms of the Vanilla configuration in Section 5 and the one using RMSNorms instead of LNs (“RMSNorm” in figures) with 350M parameters. We trained them with $\mathrm{lr}=5.0\times 10^{-4}$ as in Section 5131313We tried to train them with $\mathrm{lr}=1.0\times 10^{-3}$ but RMSNorm exploded.. As shown in these figures, RMSNorm faced loss and gradient spikes in a similar location to the ones of Vanilla. These empirical results also indicate that the RMSNorms have the same problem as LNs regarding the instability.
(a) Loss curves in validation data.
(b) Gradient norms.
Figure 10: Loss values and gradient norms of Vanilla and RMSNorm.
Appendix E Scaling Embeddings with Larger Value
Equations (15) and (20) indicate that we can stabilize the LLM pre-training by adjusting the standard deviation of the shortcut to a large value. In fact, our experimental results show that we can stabilize the LLM pre-training by making the standard deviation of each embedding close to $1$. To investigate how about a larger value, we conducted experiments with making the standard deviation of each embedding close to 5 and 50.
Figure 11 shows loss curves in validation data for 350M and 1.7B parameters in each situation. This figure indicates that although all settings prevented the loss spike problem, the larger standard deviation than $1$ degraded the performance. Therefore, it is unnecessary to scale the standard deviation of each embedding with a larger value than $1$ to prevent the performance degradation.
(a) 350M parameters.
(b) 1.7B parameters.
Figure 11: Loss curves in validation data when we scale the standard deviation of embeddings with larger than $1$.
Appendix F Distributions of Sub-Layer Outputs
Figure 12 shows output distributions of each sub-layer for each layer at the initialization. This figure indicates that each sub-layer output is close to the normal distribution in various configurations. Therefore, the assumption in this study, which is that the vector $x$ at each layer follows the normal distribution, is reasonable.
(a) self-attention of 350M parameters.
(b) self-attention of 1.7B parameters.
(c) self-attention of 13B parameters.
(d) FFN of 350M parameters.
(e) FFN of 1.7B parameters.
(f) FFN of 13B parameters.
Figure 12: Output distributions of each sub-layer.
Appendix G Discussion on ReLU in FFN
We consider the case where we use the ReLU function as $\mathcal{F}$ instead of the identity function. Because we assume that parameters and the input vector at each layer follow the normal distribution, the internal layer also follows the normal distribution. Therefore, each element of the FFN internal layer is a negative value with half probability. In this case, we can regard that the ReLU function cuts the elements of the internal layer by half. Thus, we replace $W_{1}\in\mathbb{R}^{d_{\mathrm{ffn}}\times d}$ and $W_{2}\in\mathbb{R}^{d\times d_{\mathrm{ffn}}}$ with $W_{1}\in\mathbb{R}^{\frac{d_{\mathrm{ffn}}}{2}\times d}$ and $W_{2}\in\mathbb{R}^{d\times\frac{d_{\mathrm{ffn}}}{2}}$ in the discussion in Section 3.1 when we use the ReLU function as $\mathcal{F}$.
Appendix H Relation between Input Length and the Standard Deviation of Self-Attention
We explain that the standard deviation of the self-attention layer becomes small as the input length is long. Because we assume that parameters and the input vector at each layer follow the normal distribution, the expectation of each element of $(W_{Qi}\ x)^{\mathrm{T}}(W_{Ki}\ X)$ is 00. Therefore, the expectation after the softmax function is $\frac{1}{L}$ where $L$ is the length of the input sequence. Thus, the long input sequence decreases the standard deviation of the self-attention layer.
To simplify, we consider the case where the number of self-attention heads is $1$. In this case, we can obtain the variance of each calculation with the following equation.
$\mathrm{var}(W_{O}(x))$ $=\mathrm{var}(W_{O})\mathrm{var}(x)\ d$ (23)
$\mathrm{var}(W_{V}(x))$ $=\mathrm{var}(\mathrm{W_{V}})\mathrm{var}(x)\ d$ (24)
where $\mathrm{var}$ represents the variance of the matrix/vector. Thus, the variance of the self-attention layer, $\mathrm{var}(\mathrm{Attn}(x))$, is as follows:
$\mathrm{var}(\mathrm{Attn}(x))$ $=\mathrm{var}(W_{O})\ d\sum^{L}\frac{\mathrm{var}(\mathrm{W_{V}})\mathrm{var}(x)\ d}{L^{2}}$ (25)
$=\frac{\mathrm{var}(W_{O})\mathrm{var}(\mathrm{W_{V}})\mathrm{var}(x)d^{2}}{L}$ (26)
Appendix I Details on Jacobian Matrix of Self-Attention
We can represent $\mathrm{concat}(\mathrm{head}_{1}(x),...,\mathrm{head}_{h}(x))$ with the summation of the matrix multiplications as follows:
$\mathrm{concat}(\mathrm{head}_{1}(x),...,\mathrm{head}_{h}(x))=\sum_{i=1}^{h}\mathrm{head}_{i}W_{i}$ (27)
where $W_{i}\in\mathbb{R}^{d_{\mathrm{head}}\times d}$ whose corresponding element is $1$ and the others are 00. Let $J^{i}$ be the Jacobian of the $\mathrm{head}_{i}(x)$, we can represent $J^{Z}$ in Section 3.2 as follows:
$J^{Z}=\sum_{i=1}^{h}J^{i}W_{i}$ (28)
In addition, the self-attention consists of the interaction among inputs and outputs of each position in the sequence. Thus, we add indices to Jacobians to represent the positions of the input and output. Let $x_{j}$ be the input of the position $j$, and $z_{k}^{i}$ be the $i$-th head of the output position $k$, and $J_{kj}^{i}=\frac{\partial z_{k}^{i}}{\partial x_{j}}$. Because $J^{Z}$ can be regarded as the Jacobian of the input of the position $j$, we can convert Equation (29) into the following Equation:
$J^{Z}=\sum_{k=1}^{L}\sum_{i=1}^{h}J_{kj}^{i}W_{i}$ (29)
where $L$ is the length of the input and output sequences. Therefore, we compute $J_{kj}^{i}$ to obtain $J^{Z}$ in Section 3.2.
We can obtain a head of the output position $k$, i.e., $z_{k}$, as follows [^14] To simplify the equations, we omit the index $i$ to represent $i$-th head from the head and parameters.
$z_{k}=\sum_{l=1}^{L}A_{kl}v_{l}$ (30)
where $A_{kl}$ is the $l$-th element of the attention vector, $\mathrm{softmax}\left(\frac{(W_{Q}\ x_{k})^{\mathrm{T}}(W_{K}\ X)}{\sqrt{d_{\mathrm{head}}}}\right)$ and $v_{l}$ is $W_{V}x_{l}$. Therefore, to obtain $J_{kj}$, we differentiate $z_{k}$ with respect to $x_{j}$ as:
$J_{kj}=\frac{\partial z_{k}}{\partial x_{j}}=\sum_{l=1}^{L}\left(\frac{\partial A_{kl}}{\partial x_{j}}{v_{l}}^{\mathrm{T}}+A_{kl}\frac{\partial v_{l}}{\partial x_{j}}\right)$ (31)
$\frac{\partial v_{l}}{\partial x_{j}}$ $=W_{V}\delta_{lj}$ (32)
$\delta_{lj}$ =$=\begin{cases}1\ \ \text{if $l=j$}\\ 0\ \ \text{otherwise}\end{cases}$ (33)
Thus,
$\frac{\partial z_{k}}{\partial x_{j}}=\sum_{l=1}^{L}\left(\frac{\partial A_{kl}}{\partial x_{j}}{v_{l}}^{\mathrm{T}}\right)+A_{kj}W_{V}$ (34)
Here, we assume that the attention vector is uniform. In this assumption, since $A_{kj}=\frac{1}{L}$, we can obtain the spectral norm of the second term for Equation (34) as $\|A_{kj}W_{V}\|_{2}\approx\frac{\sigma_{V}}{L}(\sqrt{d}+\sqrt{d_{\mathrm{head}}})$, where $\sigma_{V}$ is the standard deviation of $W_{V}$. To calculate the first term, we compute $\frac{\partial A_{kl}}{\partial x_{j}}$.
$\frac{\partial A_{kl}}{\partial x_{j}}$ $=A_{kl}\left(\frac{\partial S_{kl}}{\partial x_{j}}-\sum_{m=1}^{L}A_{km}\frac{\partial S_{km}}{\partial x_{j}}\right)$ (35)
$\frac{\partial S_{kl}}{\partial x_{j}}$ $=\frac{1}{\sqrt{d_{\mathrm{head}}}}\left(W_{Q}^{\mathrm{T}}W_{K}x_{l}\delta_{kj}+W_{K}^{\mathrm{T}}W_{Q}x_{k}\delta_{lj}\right)$ (36)
Let $D_{l}$ be $W_{Q}^{\mathrm{T}}W_{K}x_{l}$ and $E_{k}$ be $W_{K}^{\mathrm{T}}W_{Q}x_{k}$. Then,
$\frac{\partial S_{kl}}{\partial x_{j}}$ $=\frac{1}{\sqrt{d_{\mathrm{head}}}}\left(D_{l}\delta_{kj}+E_{k}\delta_{lj}\right)$ (37)
Therefore,
$\frac{\partial A_{kl}}{\partial x_{j}}$ $=\frac{A_{kl}}{\sqrt{d_{\mathrm{head}}}}\left((D_{l}\delta_{kj}+E_{k}\delta_{lj})-\sum_{m=1}^{L}A_{km}(D_{m}\delta_{kj}+E_{k}\delta_{mj})\right)$ (38)
$=\frac{A_{kl}}{\sqrt{d_{\mathrm{head}}}}\left(\delta_{lj}E_{k}-A_{kj}E_{k}+\delta_{kj}\left(D_{l}-\sum_{m=1}^{L}A_{km}D_{m}\right)\right)$ (39)
We assign Equation (39) to the first term of Equation (34) and use the assumption $A_{kj}=\frac{1}{L}$:
$\sum_{l=1}^{L}\left(\frac{\partial A_{kl}}{\partial x_{j}}{v_{l}}^{\mathrm{T}}\right)$ $=\sum_{l=1}^{L}\left(\frac{A_{kl}}{\sqrt{d_{\mathrm{head}}}}\left(\delta_{lj}E_{k}-A_{kj}E_{k}+\delta_{kj}\left(D_{l}-\sum_{m=1}^{L}A_{km}D_{m}\right){v_{l}}^{\mathrm{T}}\right)\right)$ (40)
$=\frac{1}{L\sqrt{d_{\mathrm{head}}}}\sum_{l=1}^{L}\left(\delta_{lj}E_{k}-\frac{1}{L}E_{k}+\delta_{kj}\left(D_{l}-\sum_{m=1}^{L}\frac{1}{L}D_{m}\right){v_{l}}^{\mathrm{T}}\right)$ (41)
$=\frac{1}{L\sqrt{d_{\mathrm{head}}}}\left(\sum_{l=1}^{L}\left(\left(\delta_{lj}E_{k}+\delta_{kj}D_{l}\right){v_{l}}^{\mathrm{T}}\right)+\left(\sum_{l=1}^{L}\left(-\frac{1}{L}E_{k}-\sum_{m=1}^{L}\frac{1}{L}D_{m}\right){v_{l}}^{\mathrm{T}}\right)\right)$ (42)
$=\frac{1}{L\sqrt{d_{\mathrm{head}}}}\left(E_{k}{v_{j}}^{\mathrm{T}}+\sum_{l=1}^{L}\left(\delta_{kj}D_{l}{v_{l}}^{\mathrm{T}}\right)-\left(E_{k}+\sum_{m=1}^{L}D_{m}\right){v_{l}}^{\mathrm{T}}\right)$ (43)
$=\frac{1}{L\sqrt{d_{\mathrm{head}}}}\left(E_{k}{v_{j}}^{\mathrm{T}}-E_{k}{v_{l}}^{\mathrm{T}}+\delta_{kj}\sum_{l=1}^{L}\left(D_{l}{v_{l}}^{\mathrm{T}}\right)-\left(\sum_{m=1}^{L}D_{m}\right){v_{l}}^{\mathrm{T}}\right)$ (44)
Based on the assumption that parameters and the vector at each layer follow the normal distribution, we assume that the mean of $D_{l}$, $E_{k}$, and $v_{l}$ is 00, and thus, we can obtain their norms from their variances. In addition, we assume that the standard deviation of $W_{Q}$, $W_{K}$ and $W_{V}$ is $\sigma$. Then, $\mathrm{var}(D_{l})=\mathrm{var}(E_{k})=dd_{\mathrm{head}}\sigma^{4}$, $\mathrm{var}(\sum_{l}^{L}D_{l})=Ldd_{\mathrm{head}}\sigma^{4}$, and $\mathrm{var}(v_{l})=d\sigma^{2}$. Thus, $\|E_{k}v_{j}^{\mathrm{T}}\|_{2}\approx\|E_{k}v_{l}^{\mathrm{T}}\|_{2}\approx\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}^{2}}$, $\|\sum_{l=1}^{L}(D_{l}v_{l}^{\mathrm{T}})\|_{2}\approx\|(\sum_{m=1}^{L}D_{m})v_{l}^{\mathrm{T}}\|_{2}\approx\sigma^{3}\sqrt{Ld^{3}d_{\mathrm{head}}^{2}}$. Therefore,
$\Bigg{\|}\sum_{l=1}^{L}\left(\frac{\partial A_{kl}}{\partial x_{j}}{v_{l}}^{\mathrm{T}}\right)\Bigg{\|}_{2}$
$\leq\frac{1}{L\sqrt{d_{\mathrm{head}}}}\left(\Big{\|}E_{k}{v_{j}}^{\mathrm{T}}\Big{\|}_{2}+\Big{\|}E_{k}{v_{l}}^{\mathrm{T}}\Big{\|}_{2}+\delta_{kj}\Bigg{\|}\sum_{l=1}^{L}\left(D_{l}{v_{l}}^{\mathrm{T}}\right)\Bigg{\|}_{2}+\Bigg{\|}\left(\sum_{m=1}^{L}D_{m}\right){v_{l}}^{\mathrm{T}}\Bigg{\|}_{2}\right)$ (45)
$\approx\frac{1}{L\sqrt{d_{\mathrm{head}}}}\left(2\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}^{2}}+(\delta_{kj}+1)\sigma^{3}\sqrt{Ld^{3}d_{\mathrm{head}}^{2}}\right)$ (46)
$=\frac{1}{\sqrt{L}}\left(\delta_{kj}+1+\frac{2}{\sqrt{L}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}$ (47)
Based on this Equation, we can compute the upper bound of $\|J_{kj}\|_{2}$:
$\|J_{kj}\|_{2}$ $=\Bigg{\|}\sum_{l=1}^{L}\left(\frac{\partial A_{kl}}{\partial x_{j}}{v_{l}}^{\mathrm{T}}\right)+A_{kj}W_{V}\Bigg{\|}_{2}$
$\leq\Bigg{\|}\sum_{l=1}^{L}\left(\frac{\partial A_{kl}}{\partial x_{j}}{v_{l}}^{\mathrm{T}}\right)\Bigg{\|}_{2}+\|A_{kj}W_{V}\|_{2}$ (48)
$\leq\frac{1}{\sqrt{L}}\left(\delta_{kj}+1+\frac{2}{\sqrt{L}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\frac{\sigma}{L}(\sqrt{d}+\sqrt{d_{\mathrm{head}}})$ (49)
Moreover, we can compute the upper bound of $\|J^{Z}\|_{2}$ from Equation (29) as follows:
$\|J^{Z}\|_{2}$ $=\bigg{\|}\sum_{k=1}^{L}\sum_{i=1}^{h}J_{kj}^{i}W_{i}\bigg{\|}_{2}$ (50)
$\leq\sum_{k=1}^{L}\sum_{i=1}^{h}\|J_{kj}^{i}\|_{2}\|W_{i}\|_{2}$ (51)
$\approx h\left(\frac{1}{\sqrt{L}}\left(1+L+\frac{2L}{\sqrt{L}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\frac{L\sigma}{L}(\sqrt{d}+\sqrt{d_{\mathrm{head}}})\right)$ (52)
$=h\left(\left(\sqrt{L}+2+\frac{1}{\sqrt{L}}\right)\sigma^{3}\sqrt{d^{3}d_{\mathrm{head}}}+\sigma(\sqrt{d}+\sqrt{d_{\mathrm{head}}})\right)$ (53)
Appendix J Comparison with Post-LN Transformer